given that the dominant isotope mass for copper metal is 62.94 amu (g/mol) how many atoms of copper metal are in a pure copper coin that weighs 2.15g
1
Expert's answer
2015-12-28T09:52:04-0500
n(Cu) = 2.15 / 62.94 = 0.0342 (mol);
N(Cu) = 0.0342 * 6.02 *1023 = 2.06 *1022

Answer: 2.06 *1022 atoms